OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

g(x) = x^2 +2 g(0) = 0^2 + 2 g(-1) = (-1)^2 +2 g(100) = (100)^2 +2

OpenStudy (anonymous):

so g(3) = ?